# DPCL-Diff: The Temporal Knowledge Graph Reasoning based on Graph Node Diffusion Model with Dual-Domain Periodic Contrastive Learning

### v1

@@ 4.1 Experimental Setup
Baselines. DPCL-Diff is compared with both static and temporal KG reasoning models. Static KG reasoning models, which do not consider temporal information, include ComplEx Trouillon et al. (2016), R-GCN Schlichtkrull et al. (2018) and ConvE Dettmers et al. (2018). In contrast, temporal KG reasoning models are designed to process temporal information and capture evolving patterns in TKG, such as RE-Net Jin et al. (2019), xERTE Z et al. (2020), CyGNet Zhu et al. (2021), EvoKG Park et al. (2022), RPC Liang et al. (2023), CENET Xu et al. (2023), DHE-TKG Liu et al. (2024b), RLGNet Lv et al. (2024), and DiffuTKG Cai et al. (2024).
Implementation Details. For all datasets, the initial embedding dimension is configured as 200. We set the batch size to 64, the embedding dimension to 200, and the learning rate to 0.001, and we used the Adam optimizer. The training calendar element limit for L is 30, and the comparison learning calendar element limit for the second stage is 20. For baseline settings, we use their recommended configuration. All experiments were implemented on servers with RTX 3090 GPUs equipped with Intel Xeon Gold 6330 CPUs.
Method YAGO ICEWS18
MRR Hits@1 Hits@5 Hits@10 MRR Hits@1 Hits@5 Hits@10
w/o- GNDif 81.30 80.64 81.42 82.26 44.93 38.22 47.53 57.08
w/o- DPCL 84.21 84.19 84.23 84.34 46.51 39.92 48.92 59.67
DPCL-Diff 84.45 84.23 84.65 84.75 47.02 40.03 49.93 60.66
@@
4.2 Main Results
We compared the experimental results in Table 1 and Table 2, which provides the MRR and Hits@1/3/10 metrics of DPCL-Diff on four datasets. DPCL-Diff outperforms 13 other comparison models in most cases, particularly excelling on ICEWS14, where it achieves a 29.54% improvement in Hits@1 over our baseline model CENET. This significant gain is mainly due to effective data generation techniques, enabling the model to handle unseen events more effectively. ICEWS14 has a high proportion of new events(about 30% Xu et al. (2023)) and contains the most significant volume of data among the four datasets. DPCL-Diff generates substantial-high-quality data through GNDiff, enhancing inference on these events and improving overall performance. / Moreover, while DiffuTKG was the first to introduce diffusion processes in TKG inference, DPCL-Diff further leverages a pre-trained language model (PLM) in the denoising process, capturing contextual information more effectively and aligning data generation closer to actual event distributions. This is particularly advantageous in scenarios with sparse interaction trajectories, leading to superior predictive accuracy and allowing DPCL-Diff to outperform DiffuTKG across key metrics, especially in datasets with a high proportion of new events./ For the YAGO and WIKI datasets, DPCL-Diff also outperforms other algorithms on most metrics, except for Hits@10 on YAGO, which is slightly lower than RLGNet but significantly higher in Hits@1. DPCL-Diff ability to distinguish similar periodic event entities improves prediction accuracy. Although YAGO and WIKI have a lower proportion of new events (about 10% Xu et al. (2023)) and a higher percentage of periodic events, which limits GNDiff performance, DPCL-Given that the GNDiff parameters used in this experiment are relatively low, with the batch size set to 64, these settings may influence the results. We believe the performance could be significantly improved if the batch size is increased to 512 or 1024.
4.3 Ablation Study
@@
To investigate the effectiveness of the modules in DPCL-Diff, we verified the effectiveness of GNDiff and DPCL through ablation experiments. The ablation results are shown in Table 4. Here, "w/o-GN Diff" denotes the removal of the graph node diffusion model, and "w/o-DPCL" denotes the removal of the Poincaré mapping, which maps both positive and negative samples into the Euclidean space for contrast learning. These results illustrate the effectiveness of our proposed GNDiff and DPCL. The performance degradation is particularly noticeable after the removal of GNDiff. This demonstrates that generating a large number of high-quality data related to the predicted event through GNDiff significantly affects event prediction. DPCL leverages Poincaré space properties to distinguish similar periodic events better, enhancing inference accuracy.
@@
4.4 Dual Domain Mapping Strategy Analysis
In this study, we separately examined the effects of different spatial mapping strategies on the learning effect of DPCL. Table 4 shows the results of the experiments on the YAGO dataset without the GNDiff. Euc/Hyp denotes that periodic and non-periodic event entities are mapped to Euclidean and Poincaré spaces respectively for comparative learning, Hyp/ Euc denotes that the two types of entities are mapped to Poincaré and Euclidean space, Hyp/Hyp denotes that both types of entities are mapped to Poincaré space, and Euc / Euc denotes that the two types of entities are mapped to Euclidean space respectively. From the experimental results, it can be seen that the Hyp/Euc method is the most effective, while the Euc/Euc method is the least effective. This suggests that mapping periodic and non-periodic event entities to separate spaces can improve the modeling of comparative learning. Specifically, the geometric properties of the Poincaré space can effectively distinguish entities whose feature representations are too close to each other in the Euclidean space.
5 Conclusion
@@
6 Limitations
This study does not incorporate adaptive embedding strategies to distinguish between periodic and non-periodic events. Consequently, our approach lacks the flexibility to dynamically adjust to different types of temporal knowledge graph data. This limitation could potentially impact the model’s effectiveness when applied to datasets with varying temporal characteristics.
Appendices
7 Extended Experiments
7.1 Dual Domain embedded visual analytics
In this section, we investigate the visualization of entities in DPCL with bi-domain embedding. As shown in Fig. 2, we visualize two sets of entity pairs—(Germany, Berlin) and (Germany, Munich)—with the same relation, mapping them to Euclidean and Poincaré space, respectively. In Euclidean space, due to the similar Euclidean distance between Berlin and Munich to Germany, it is difficult to distinguish between these two entities effectively. However, in Poincaré space, Germany is significantly farther from Munich, bringing Germany closer to Berlin and making it easier to derive the correct inference. By leveraging the spatial properties of Poincaré space, we can more effectively distinguish similar periodic event entities and improve the overall prediction accuracy of the model.

### v2

@@ Introduction
Method YAGO
MRR Hits@1 Hits@3 Hits@10
ComplEx(2016) 61.29 54.88 62.28 69.42
R-GCN(2018) 41.3 32.56 44.44 ——
ConvE(2018) 62.32 56.19 63.97 70.44
RE-Net(2019) 65.16 63.29 65.63 68.08
xERTE(2020) 58.75 58.46 58.85 60.48
CyGNet(2021) 63.47 64.26 65.71 68.95
EvoKG(2022) 55.11 54.37 81.38 92.73
CENET(2023) 79.64 78.65 79.91 81.42
DHE-TKG(2024) 62.93 71.00 82.72
RLGNet(2024) 80.17 76.52 83.57 84.96
DPCL-Diff 84.45 84.23 84.65 84.75
Improve(%) 6.04 $\uparrow$ 7.09 $\uparrow$ 5.86 $\uparrow$ 4.09 $\uparrow$
Table 3: Experimental results of temporal link prediction on YAGO.
Baselines. DPCL-Diff is compared with both static and temporal KG reasoning models. Static KG reasoning models, which do not consider temporal information, include ComplEx (Trouillon et al. 2016), R-GCN (Schlichtkrull et al. 2018) and ConvE (Dettmers et al. 2018). In contrast, temporal KG reasoning models are designed to process temporal information and capture evolving patterns in TKG, such as RE-Net (Jin et al. 2019), xERTE (Z et al. 2020), CyGNet (Zhu et al. 2021), EvoKG (Park et al. 2022), RPC (Liang et al. 2023), CENET (Xu et al. 2023), DHE-TKG (Liu et al. 2024b), RLGNet (Lv et al. 2024), and HisRES (Zhang et al. 2024).
Method ICEWS18 YAGO
MRR Hits@1 Hits@3 Hits@10 MRR Hits@1 Hits@3 Hits@10
w/o- GNDif 44.93 38.22 47.53 57.08 81.30 80.64 81.42 82.26
w/o- DPCL 46.51 39.92 48.92 59.67 84.21 84.19 84.23 84.34
DPCL-Diff 47.02 40.03 49.93 60.66 84.45 84.23 84.65 84.75
@@
Method ICEWS14(new events) ICEWS18(new events) WIKI(new events)
MRR Hits@1 Hits@3 Hits@10 MRR Hits@1 Hits@3 Hits@10 MRR Hits@1 Hits@3 Hits@10
GNDiff 47.36 41.56 49.04 58.72 37.46 31.64 39.71 48.42 62.51 61.97 62.77 63.37
DPCL 42.37 36.41 44.78 53.03 33.93 29.47 35.21 42.18 57.16 56.81 57.28 57.58
Table 5: Performance of Graph Diffusion Models on New Events in ICEWS14 and ICEWS18 and WIKI.
Method YAGO(new events)
MRR Hits@1 Hits@3 Hits@10
GNDiff 79.49 79.44 79.48 79.54
DPCL 71.88 71.78 71.89 71.98
Table 6: Performance of Graph Diffusion Models on New Events in YAGO.
New Events Dataset #Train #Valid #Test
ICEWS18 221,000 31,407 33,211
ICEWS14 275,762 35,744 34,641
WIKI 113,629 30,317 28,811
YAGO 44,949 10,704 10,985
Table 7: Statistics of New Events Datasets.
Implementation Details. For all datasets, DPCL uses an embedding dimension of 200, while GNDiff uses 128. We set the batch size to 64, the learning rate to 0.001, and used the Adam optimizer. The training calendar element limit for $L$ is 30, and the comparison learning calendar element limit for the second stage is 20. The value of the hyperparameter $\alpha$ is set to 0.2, and $\lambda$ is set to 2. Additionally, $\tau$ is the temperature parameter set to 0.1 in the supervised contrastive loss $\mathcal{L}_{sup}$. For baseline settings, we use their recommended configuration. All experiments were implemented on servers with RTX 3090 GPUs equipped with Intel Xeon Gold 6330 CPUs.
Main Results
We compared the experimental results in Table 1 and Table 3, presenting the MRR and Hits@1/3/10 metrics of DPCL-Diff across four datasets. DPCL-Diff consistently outperforms 12 baseline models, with a notable 30.37%(Hits@1) improvement over CENET on ICEWS14. This substantial gain is not solely attributable to the effective data generation of GNDiff for handling unseen events but also to its ability to generalize on periodic events. ICEWS14 high proportion of new events and large data volume amplify GNDiff impact, while the pre-trained language model (PLM) further enhances contextual representation. Additionally, DPCL-Diff’s ability to model structural and semantic relationships across events improves performance on both new and periodic events, with the sensitivity of the Hits@1 metric to ranking adjustments further emphasizing the improvement.
On the YAGO and WIKI datasets, DPCL-Diff also achieves strong overall performance, with significantly higher Hits@1 compared to other models, even though Hits@10 on YAGO is slightly lower than RLGNet. The model’s ability to capture relationships between periodic events enhances prediction accuracy despite the lower proportion of new events and a higher proportion of periodic events, which limit GNDiff’s direct impact. Furthermore, the experiments were conducted with a relatively small batch size of 64, potentially constraining the model’s performance. Increasing the batch size to 512 or 1024 could further enhance its capacity to capture contextual and relational information, especially in datasets with more complex event patterns.
Ablation Study
@@
To investigate the effectiveness of the modules in DPCL-Diff, we verified the effectiveness of GNDiff and DPCL through ablation experiments. The ablation results are shown in Table 4. Here, ”w/o-GN Diff” denotes the removal of the graph node diffusion model, and ”w/o-DPCL” denotes the removal of the Poincaré mapping, which maps both positive and negative samples into the Euclidean space for contrast learning. These results illustrate the effectiveness of our proposed GNDiff and DPCL. The performance degradation is particularly noticeable after the removal of GNDiff. This demonstrates that generating a large number of high-quality data related to the predicted event through GNDiff significantly affects event prediction. DPCL leverages the properties of Poincaré space to better capture the relationships between periodic entities and distinguish similar periodic events, thereby enhancing inference accuracy.
Effectiveness Analysis of Graph Diffusion in New Events
To evaluate the effectiveness of graph diffusion models on new events, we curated datasets from four sources by selecting the first occurrences of distinct events, thereby focusing exclusively on new data. We extracted datasets based on the first appearance of all events that occurred, ensuring that only new and unique occurrences were included, while subsequent appearances of the same events were excluded. Detailed dataset statistics are provided in Table 7, which supports our methodological approach. The performance comparisons, detailed in Table 5 and Table 6, show that the GNDiff model consistently outperforms the DPCL model across all metrics and datasets. This robust performance highlights the model’s capability to adeptly handle dynamic, unseen data, proving GNDiff’s exceptional utility in environments that require the continuous integration and processing of new information.
@@
Dual Domain Mapping Strategy Analysis
In this study, we examined the effects of different spatial mapping strategies on DPCL’s learning performance. Table 8 presents the results on the YAGO dataset without GNDiff. Euc/Hyp denotes that periodic and non-periodic event entities are mapped to Euclidean and Poincaré spaces, respectively, for comparative learning, Hyp/ Euc denotes that the two types of entities are mapped to Poincaré and Euclidean space, Hyp/Hyp denotes that both types of entities are mapped to Poincaré space. Euc / Euc denotes that the two types of entities are mapped to Euclidean space, respectively. The experimental results show that the Hyp/Euc method is the most effective, while the Euc/Euc method is the least effective. This suggests that mapping periodic and non-periodic event entities to separate spaces can improve the modelling of comparative learning. Specifically, compared to Euclidean space, the geometric properties of the Poincaré space allow for a more accurate capture of the relationships between periodic entities and distinguish similar ones, leading to better link prediction results.
Dual Domain Embedded Visual Analytics
In this section, we visualize entity pairs—(Germany, Berlin) and (Germany, Munich)—with the same relation, embedded in both Euclidean and Poincaré spaces, as shown in Fig. 2. In Euclidean space, the relationships between Germany and the two cities are not well differentiated. In contrast, Poincaré space provides a clearer distinction, better reflecting the relationships between the entities. By leveraging the geometric properties of Poincaré space, we improve the model’s ability to capture entity relationships, better distinguish similar entities, thereby enhancing prediction accuracy.
Graph Node Diffusion Generation Process
To address the issue of insufficient interaction traces for new events on the timeline, GNDiff generates data samples by simulating new events using a diffusion mechanism. Table 9 illustrates this process for GNDiff. Since graph node data differs from text data, we first obtain node embedding vectors and then apply the diffusion mechanism. Each node is transformed into a 128-dimensional vector, but for simplicity, we use a 4-dimensional vector to demonstrate the process.
The process begins with a fully masked state and progressively reveals information about entities and relationships. Initially (t = 0), all information is masked. For instance, even though entity 1 (United States) and the relation (sanctioned) are known, they are masked, requiring the model to predict entity 2. By step three (t = 3), the model predicts entity 2 based on frequency information, generating a vector [0.23, 0.37, 0.52, ?]. In step four (t = 4), the model adjusts this vector to [0.18, 0.29, 0.71, ?]. Finally, the model generates the complete vector [0.18, 0.29, 0.71, 0.86], mapping it to the entity ”Russia,” thus completing the generation process. This iterative adjustment enhances the quality of samples, improving the model’s accuracy and robustness in predicting new events.
Hyper-parameter Analysis
DPCL-Diff involves two key hyperparameters, $\alpha$ and $\lambda$, whose impact on performance is analyzed for the ICEWS14 and YAGO datasets in Fig. 3. $\alpha$ balances $L_{\text{diff}}$ and $L_{\text{dpcl}}$. On ICEWS14, $\alpha=0.3$ enhances temporal interaction modeling with minimal sensitivity to changes. For YAGO, a lower $\alpha$ is more effective, highlighting the importance of dual-domain contrastive learning for periodic events. For $\lambda$, after fixing $\alpha$, $\lambda=4$ optimizes performance on ICEWS14 by balancing temporal variations and preventing overfitting. On YAGO, progressively increasing $\lambda$ improves performance, handling the dataset’s complex structure. These results demonstrate DPCL-Diff’s robustness and ability to maintain high performance with stable hyperparameter settings, effectively balancing diffusion and contrastive learning across datasets.
step Entity 1 Relation Entity 2
t = 0 [MASK] [MASK] [MASK]
t = 1 United States [MASK] [MASK]
t = 2 United States sanctioned [MASK]
t = 3 United States sanctioned [0.23, 0.37, 0.52, ?]
t = 4 United States sanctioned [0.18, 0.29, 0.71, ?]
t = 4 United States sanctioned [0.18, 0.29, 0.71, 0.86]
t = 5 United States sanctioned Russia
Table 9: Example of graphical node diffusion model generation process
Figure 3: Results of hyper-parameters $\alpha$ and $\lambda$ of DPCL-Diff on ICEWS14 and YAGO.
Conclusion
@@
Acknowledgments
This work was supported by the National Natural Science Foundation of China (No. 62402307, No. U23B2021).